40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1692     Entry time: Fri Jul 29 16:52:06 2016
Author: awade and Antonio 
Type: DailyProgress 
Category: BEAT 
Subject: Beat note signal 

We have a beat-note.

For the last few days we have been able to bring the two lasers to within the 125 MHz bandwidth of the Newfocus 1811 detector at the output of the vacuum tank. By tuning the length of the North cavity (turning the temperature up) we were able to also bring the resonance of the cavities to within this bandwidth.  

Using the beat note detector setup we assembled at the start of the laser path (PSL_Lab/1689), we can see that the frequency offset of the north and south cavity resonance settled at 70 MHz for a north heater voltage set point of 11.5 V (74 mA). We still don't know what temperature that is because we have no thermistor monitor setup yet. However, it seems that with enough settling time it is stable enough for these initial diagnostic tests.

However, until today we had trouble finding a beat note after the two ref cavities.  We checked polarization, co-alignment of the beams as well as switching the detector in and out with the NF-1611 (1 GHz) detector borrowed from the 40-m.  It turns out there are some neutral density filters installed right before the final focusing lens in the post refcav beatnote setup. When I checked the specs for the NF-1811 it looked like we were well below the RF saturation point with a lot of room to increase power.

For the NF-1811 saturation power @ 1550 nm is 55 uW. The responsively at 1064 nm looks like about 0.6 of this.  So saturation at our wavelength is about 91.7 uW. I measured the total power after the combining beamsplitter, it was 227 uW, so with 13 dB worth of ND filters this was brought down to about 11.3 uW.  We first switched the ND filters (ND1.0+ND0.3) for a single ND0.5 and a beat note could be see 3 dB above the dark noise (75 dBm) of the detector. We then removed the ND filter completely at got a beat note (at 70 MHz) that had a clearance of 17 dB above dark noise.  Not sure if this is enough to make a 'good' measurement but at least we can now see something at the combined transmission of the cavities. The damage threshold is of course way above this.

There is quite a bit of work to be done improving on this and getting the PLL loop setup.  We also need to start work on getting a calibration together. The PDH loops also need some optimizing.

 

On another side note:

We can see the 14.75 MHz  side bands at the beat note output (and one 2x harmonic of this). I'm not sure if this is ok as I would have through the ref cavities would filter this out and I'm not sure if maybe HOM or other would let this to pass the cavity. 

ELOG V3.1.3-